OpenStudy (anonymous):

what does this mean? reduce the fraction by finding the common factor in the numerator and in the denominator and dividing by the common factor 72 --- 96

OpenStudy (bahrom7893):

i think common factor is 24

OpenStudy (bahrom7893):

72/96 = (24*3)/(24*4) = 3/4

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

thank you